# Akaki Mamageishvili Offchain Labs Benny Sudakov ETH Zurich

### v2

@@ 3.2 PoS Savings
Suppose a PoS protocol centered at any service $s\in S$, having security parameter $\alpha_{s}$ and allocated $\sigma_{v}^{s}$ stakes for any $v\in N_{G}(s)$ is safe, by the same definition as above: no attacking set of validators find it profitable to attack service $s$. Let the partition of the stake vector $\sigma$ be denoted by $\sigma^{S}$. Consider the total stake for any validator $v\in V$ defined as $\sigma_{v}=\sum_{s\in N_{G}(v)}\sigma_{v}^{s}$. That is, we aggregate all stakes that the validator has staked across all PoS protocols in which it participates. We ask a similar question to the previous section:
@@
How much stake do we need to add to the aggregate validityator stakes, $\sigma_{v}$, so that the resulting restaking graph is secure?
The security of the restaking graph is considered by the original definition 2. More formally, for each restaking graph, we define:
@@
$PoSS(G)=\min_{\sigma^{\prime}\in D(\sigma)}\frac{T(\sigma^{\prime})-T(\sigma)}{T(\sigma)}$
First, we note the following inequality that relates the total value to capture from all services and the total staked amount of validators. It will be useful to obtain an upper bound on the PoS savings.
@@ A PoS protocol centered at $s\in S$ is secure implies that
Consider PoS protocols with $m$ services, each service $s\in S$ having value $\pi_{s}=1$, security parameter $\alpha_{s}=\frac{1}{m+1}$ and two validators, labeled $s$ and m+1{m+1} with stakes equal to $\sigma_{s}^{s}=1$ and $\sigma_{m+1}^{s}=\frac{1}{m}$. Note that all PoS protocols are secure:
  • large staked validator $s$ does not attack since it is burning stake equal to the value it obtains,
  • small staked validator m+1{m+1} can not attack the service since its stake share is equal to $\frac{\frac{1}{m}}{1+\frac{1}{m}}=\frac{1}{m+1}$, that is the same as a security parameter of the service.
We claim that the minimum additional stake required is m1m-1. In the resulting restaking protocol, the validator m+1{m+1} has stake equal to $m\frac{1}{m}=1$. It can attack all services unless one of the other validators gets an extra m1m-1 stake, in which case the claim is trivially true. This amount is enough. Consider adding stake m1m-1 to the validator m+1{m+1} so that it now controls the stake amount $m$. This validator does not attack services since it is losing at least the amount she can gain by attacking. Other validators are not able to attack their corresponding services, and it is also not profitable to attack.
Hence, $PoSS\geq\frac{m-1}{m+1}$ for any $m$, since the original total stake amount among the validators was $m+1$. Taking $m\rightarrow\infty$ gives $PoSS\geq 1$. ∎
@@
For each service $s\in S$, add to all its validators $v\in N_{G}(s)$ a stake equal to $\pi_{s}$. The resulting restaking graph is secure. For any subset of validators, $U\subseteq V$, they lose more stakes than the value they derive from the maximal set they attack, $M(U)$. The reason for this is that by our construction, at least one of the validators in $U$ received an additional stake of $\pi_{s}$, for any $s\in M(U)$. We added an extra stake of
@@
where the last inequality is derived in Lemma 1. This completes the proof of the claim of the proposition. ∎
@@
The upper bound obtained in Proposition 4 is interesting in practical cases, since the security parameters of most protocols are usually from the set of fractions $\{\frac{1}{4},\frac{1}{3},\frac{1}{2}\}$. These values come from Byzantine fault-tolerant mechanisms, used to achieve consensus on the state of the chain.
The number of unique addressed validators is high for the Ethereum chain, approaching one million, while the real number of validators can still be on the order of tens of thousands. All validators are staked for Ethereum, that is, the restaking protocol based on Ethereum has the highest degree $\max_{s}{d_{G}(s)}=n$. The number of projects that require staking services is typically low, in the order of hundreds, and it is highly unlikely that a validator would provide services for most of them, therefore, it would be expected that $\max_{v}(d_{G}(v))$ is low.
Recently, EigenLayer’s restaking project upgraded the system that only allows a unique slashing property: any staked asset can be slashed by a single service  [^4] https://blog.eigencloud.xyz/introducing-the-eigenlayer-security-model/. This makes it equivalent to validators distributing their total stake amount for separate services, and hence results into separate secure PoS protocols.
@@ 4 Conclusion
We initiate a framework to compare restaking and implied Proof-of-Stake protocols in terms of total stake requirements. We provide asymptotically matching lower and upper bounds on restaking savings, as well as a constant lower bound on the PoS savings and non-constant upper bounds on it. Our results indicate that none of the solution concepts dominates another. Closing the gap for PoS savings remains an open question. Another open question is to come up with a (generic) sufficient condition for a restaking graph to be secure, that can be checked in a polynomial time, and the condition accepts some graphs with positive restaking savings.

### v3

@@ 3.2 PoS Savings
Suppose a PoS protocol centered at any service $s\in S$, having security parameter $\alpha_{s}$ and allocated $\sigma_{v}^{s}$ stakes for any $v\in N_{G}(s)$ is safe, by the same definition as above: no attacking set of validators finds it profitable to attack service $s$. Consider the total stake for any validator $v\in V$ defined as $\sigma_{v}=\sum_{s\in N_{G}(v)}\sigma_{v}^{s}$. That is, we aggregate all stakes that the validator has staked across all PoS protocols in which it participates. We ask a similar question to the previous section:
@@
How much stake do we need to add to the aggregate validator stakes, $\sigma_{v}$, so that the resulting restaking graph is secure?
The security of the restaking graph is considered by the original definition 2. Let $U(G)$ denote the space of stake vectors for which $G$ restaking graph is secure. Then, for each restaking graph, we define:
@@
$PoSS(G)=\min_{\sigma^{\prime}\in D(\sigma)\cap U(G)}\frac{T(\sigma^{\prime})-T(\sigma)}{T(\sigma)}$
$PoSS(G)$ thus captures the inefficiency of aggregation, where interdependencies may necessitate over-collateralization to restore security. First, we note the following inequality that relates the total value to capture from all services and the total staked amount of validators. It will be useful to obtain an upper bound on the PoS savings.
@@ A PoS protocol centered at $s\in S$ is secure implies that
Consider PoS protocols with $m$ services. Each service $s\in S$ has value $\pi_{s}=1$ and a security parameter $\alpha_{s}=\frac{1}{m+1}$. There are two types of validators, labeled $s$ for each service $s\in S$ and the last validator labeled m+1{m+1}. Stakes are equal to $\sigma_{s}^{s}=1$ and $\sigma_{m+1}^{s}=\frac{1}{m}$. Note that all $m$ PoS protocols are secure:
  • A validator labeled $s$ does not attack, since it is losing a stake equal to $1$, the value it obtains from attacking a service.
  • A validator labeled m+1{m+1} can not attack the service since its stake share is equal to $\frac{\frac{1}{m}}{1+\frac{1}{m}}=\frac{1}{m+1}$. This is the same as a security parameter of the service.
We claim that the minimum additional stake required to make aggregate restaking graph secure is m1m-1. In the resulting restaking protocol, the validator m+1{m+1} has stake equal to $m\frac{1}{m}=1$. It can attack all services unless its own stake becomes $m$ or one of the other validators gets an extra m1m-1 stake, in which case the claim is trivially true. The extra amount m1m-1 is enough. Consider adding stake m1m-1 to the validator m+1{m+1} so that it now controls the stake amount $m$. This validator does not attack services since it is losing at least the amount it can gain by attacking. Other validators are not able to attack their corresponding services, and it is also not profitable to attack.
Hence, $PoSS\geq\frac{m-1}{m+1}$ for any $m$, since the original total stake amount among the validators is $m+1$. Taking $m\rightarrow\infty$ gives $PoSS\geq 1$. ∎
@@
First, we show that $PoSS(G)\leq\max_{s\in S}d_{G}(s)$, for the intuition. For each service $s\in S$, add to all its validators $v\in N_{G}(s)$ a stake equal to $\pi_{s}$. The resulting restaking graph is secure. For any subset of validators, $U\subseteq V$, they lose more cumulative stake than the value they derive from the maximal set they attack, $M(U)$, resulting into a not profitable attack. The reason for this inequality is that by our construction, for any $s\in M(U)$, at least one of the validators in $U$ received an additional stake of $\pi_{s}$. We used an extra stake of
@@
where the last inequality is derived in Lemma 1. This completes the proof of the claim. Note that in this proof, we did not use original stakes validators are endowed to. Now suppose that to each validator $v\in V$ we add a stake equal to $\max(0,\sum_{s\in N(v)}\pi_{s}-\sigma_{v})$. It is clear from the proof of the first claim that the resulting graph is secure.
In the following, we will try to upper bound the value
$E:=\sum_{v\in V}(\max(0,\sum_{s\in N_{G}(v)}\pi_{s}-\sigma_{v})).$
Let $\Delta_{v}:=\sum_{s\in N_{G}(v)}\pi_{s}-\sigma_{v}$. The set of validators can be divided into two disjoint subsets $V^{+}$ and $V^{-}$, so that $V^{+}:=\{v\in V:\Delta_{v}\geq 0\}$ and $V^{-}:=\{v\in V:\Delta_{v}<0\}$. Let $S^{+}$ denote the set $\{s\in S:N_{G}(s)\subseteq V^{+}\}$ of the services whose validators are all in $V^{+}$, and $S^{\prime}=S\setminus S^{+}$ be the set of remaining services, each having at least one validator in $V^{-}$. For $s\in S^{\prime}$, $|N_{G}(s)\cap V^{+}|\leq d_{G}(s)-1$. Then,
$E=\sum_{v\in V^{+}}\Delta_{v}=\sum_{v\in V^{+}}(\sum_{s\in N_{G}(v)}\pi_{s}-\sigma_{v})=\sum_{s\in S}\pi_{s}|N_{G}(s)\cap V^{+}|-\sum_{v\in V^{+}}\sigma_{v}.$ (3)
Next, the following chain holds:
$\sum_{s\in S}\pi_{s}|N_{G}(s)\cap V^{+}|=\sum_{s\in S^{+}}\pi_{s}|N_{G}(s)\cap V^{+}|+\sum_{s\in S^{\prime}}\pi_{s}|N_{G}(s)\cap V^{+}|\leq$
$\sum_{s\in S+}\pi_{s}d_{G}(s)+\sum_{s\in S^{\prime}}\pi_{s}(d_{G}(s)-1).$
Therefore, the extra added value to stakes can be upper bounded as
$E\leq\sum_{s\in S+}\pi_{s}d_{G}(s)+\sum_{s\in S^{\prime}}\pi_{s}(d_{G}(s)-1)-\sum_{v\in V^{+}}\sigma_{v}.$
Since each PoS protocol is secure, we have $\sum_{v\in N_{G}(s)}\sigma_{v}^{s}\geq\pi_{s}$ for all $s\in S$, which implies that:
$\sum_{s\in S^{+}}\sum_{v\in N_{G}(s)}\sigma^{s}_{v}\geq\sum_{s\in S^{+}}\pi_{s}.$
For $s\in S^{+}$, $N_{G}(s)\subseteq V^{+}$, so the left hand side of the above is:
$\sum_{v\in V^{+}}\sum_{s\in S^{+}\cap N_{G}(v)}\sigma_{v}^{s}\leq\sum_{v\in V^{+}}\sum_{s\in N_{G}(v)}\sigma_{v}^{s}=\sum_{v\in V^{+}}\sigma_{v}.$ (4)
Therefore, $\sum_{v\in V^{+}}\sigma_{v}\geq\sum_{s\in S^{+}}\pi_{s}$. Substituting in the upper bound on $E$, we get:
$E\leq\sum_{s\in S^{+}}\pi_{s}d_{G}(s)+\sum_{s\in S^{\prime}}\pi_{s}(d_{G}(s)-1)-\sum_{s\in S^{+}}\pi_{s}=\sum_{s\in S}\pi_{s}(d_{G}(s)-1).$
Finally,
$\sum_{s\in S}\pi_{s}(d_{G}(s)-1)\leq(\max_{s\in S}d_{G}(s)-1)T(\pi)\leq(\max_{s\in S}d_{G}(s)-1)T(\sigma).$
Note that the upper bound in Proposition 6 matches the lower bound limit obtained in Proposition 5, for graphs where the maximun number of validators over services is equal to two.
@@
The upper bound obtained in Proposition 4 is particularly relevant in practical cases, since the security parameters of most protocols are usually from the set of fractions $\{\frac{1}{4},\frac{1}{3},\frac{1}{2}\}$. These values come from Byzantine fault-tolerant mechanisms, used to achieve consensus on the state of the chain. In such cases, the bound simplifies to at most $2$ to $4$ times the original total stake, providing a tight estimate of the capital overhead required to achieve equivalent PoS security.
The number of unique addressed validators is high for the large-scale ecosystems such as Ethereum chain, approaching one million, while the real number of validators can still be on the order of tens of thousands. All validators are staked for Ethereum, that is, the restaking protocol based on Ethereum has the highest degree $\max_{s}{d_{G}(s)}=n$. The number of projects that require staking services is typically low, in the order of hundreds, and it is highly unlikely that a validator would provide services for most of them, therefore, it would be expected that $\max_{v}(d_{G}(v))$ remains low.
Recently, EigenLayer’s restaking project upgraded the system that only allows a unique slashing property: any staked asset can be slashed by a single service  [^6] https://blog.eigencloud.xyz/introducing-the-eigenlayer-security-model/. This restriction effectively makes it equivalent to validators partition their total stake across services, and hence results into separate secure PoS protocols. As a result, the system no longer realizes capital efficiency gains from true stake reuse, and the restaking savings collapses to zero under this constraint.
@@ 4 Conclusion
We introduce a formal framework for comparing the capital efficiency of secure restaking protocols and equivalent independent Proof-of-Stake (PoS) protocols, measured by the total stake. Our main results establish tight asymptotic bounds on the restaking savings (the relative extra capital needed to redistribute stakes from a secure restaking graph into separate secure PoS protocols): we prove matching upper and lower bounds of $\Theta(\sqrt{n})$, where $n$ is the number of validators. These bounds are achieved through greedy covering technique and explicit construction, showing that restaking can deliver meaningful capital efficiency gains in large-scale systems.
In the complementary direction, we demonstrate that aggregating secure PoS protocols into a secure restaking graph can incur a significant overhead. We provide a constant lower bound of $1$ on PoS savings (the extra stake needed post-aggregation to restore security) and non-constant upper bound equal to maximum service degree in the restaking graph.
Taken together, our results indicate that neither restaking nor independent PoS protocols dominate the other in terms of capital requirements – each has structural advantages depending on the underlying graph.
Two important open questions remain. First is closing the gap between the constant lower bound and the non-constant upper bounds on PoS savings. Second is designing a generic, polynomial-time-checkable sufficient condition for restaking graph security that admits graphs with strictly positive restaking savings.